three years ago, marc's age was 1 year more than thrice aaron's 4 years from now marc's age will be 2 years more than twice aaron's. How old is Marc?

Answer by ankor@dixie-net.com(22740)

): three years ago, marc's age was 1 year more than thrice aaron's
m-3 = 3(a-3) + 1
m - 3 = 3a - 9 + 1
m = 3a - 8 + 3
m = 3a -5
:
4 years from now marc's age will be 2 years more than twice aaron's.
m + 4 = 2(a+4) + 2
m + 4 = 2a + 8 + 2
m = 2a + 10 -4
m = 2a + 6
Replace m with (3a-5)
3a - 5 = 2a + 6
3a - 2a = 6 + 5
a = 11 yrs is Aaron's age
:
How old is Marc?
m = 3(11) - 5
m = 33 - 5
m = 28 yrs
:
Check solutions in the 2nd statement